# apndxAdditional References A Hitchhiker’s Guide to Privacy-Preserving Cryptocurrencies: A Survey on Anonymity, Confidentiality, and Auditability

### v1

@@ 4.3.3. Transaction Value Confidentiality
Two main approaches emerge from the literature to conceal transaction details: commitments (e.g., (Ben Sasson et al., 2014; Androulaki and Karame, 2014; Noether et al., 2016; Poelstra, 2016; Garman et al., 2017; Fauzi et al., 2019; Wüst et al., 2019; Gross et al., 2021; Guan et al., 2022; Wüst et al., 2022; Beer et al., 2024; Elfadul et al., 2024; Tomescu et al., 2022; Androulaki et al., 2020; Narula et al., 2018)) and public key encryption (e.g., (Kiayias et al., 2022; Sarencheh et al., 2025; Bünz et al., 2020; Ma et al., 2021)). All the approaches using commitments and most of those adopting public key encryption (e.g., (Bünz et al., 2020; Ma et al., 2021)) use ZKPs to prove correctness.
@@
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening: the opening is a random value generated and encrypted during withdrawal, and decrypted and stored at deposit time to prevent double spending. In 2013, Maxwell proposed confidential transactions121212https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness (also using range proofs). Monero adopted and extended this technique by combining ring signatures for sender’s anonymity, obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
@@
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transaction information to several maintainers who operate a distributed ledger. Transactions are encrypted using threshold ElGamal encryption (Elgamal, 1985), which requires a quorum of maintainer to cooperate to revoke confidentiality and decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. Androulaki et al. (Androulaki et al., 2020) extend Zerocoin to implement auditability: each user is assigned to an auditor at registration time, who can decrypt and inspect the user’s transactions when required. Using encryption instead of commitments allows to revoke anonymity and implement auditability and tracing. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving currencies using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define a privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, for hiding each account’s balance. Zether introduces the -Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form. We briefly mention few other works that focus on privacy in smart contracts. For example, Hawk (Saleem et al., 2024) uses MPC to split a contract into private and public parts, allowing confidential state. Enigma (Zyskind et al., 2015) uses secure enclaves to run contract code on encrypted inputs. These solutions are beyond currency transfers, but they show that account privacy often requires splitting roles or off-chain computation. Recent research projects aim for private smart contracts via ZKP frameworks, often assuming a model with private channels or truster parties (e.g., (Bowe et al., 2020)).
@@ 4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router131313https://www.torproject.org/) and I2P (Invisible Internet Project141414https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents malicious attackers from linking transactions with their source IPs. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
@@ 4.3.5. Linkability and Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user. For instance, it should be impossible to conclude that two transactions were delivered to or sent from the same user (other definitions can be found in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). Unlinkability can refer to the sender, receiver, or both. A few works break this property into two facets, internal and external. For instance, CoinJoin (Maxwell, 2013) does not preserve internal unlinkability as each participant in the group that mixes transactions can link entities within the group.
Obtaining unlinkability in public ledgers is challenging as linking between addresses in cryptocurrency systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)). A non exhaustive list of patterns and heuristics to link addresses is the following.
@@
Address reuse: Using the same address for multiple transactions make it easy to link payments to the same user151515This works because address are pseudonymous but not private.;
@@
Dust attack analysis: Small amounts of cryptocurrency are sent to many addresses to track their movements (similar to a phishing attack); if dust is later spent with other transactions, it links addresses together.
Existing countermeasures primarily aim to obscure input-output relations within transactions through mixing protocols and address reuse prevention. Mixing protocols disrupt common ownership heuristics and help mitigate change address attacks. Address-changing techniques generate new sender or receiver addresses for each transaction to prevent linkability. Additionally, ensuring network anonymity helps further obscure user identities. To counter dust attacks, the simplest defense is to ignore dust transactions, preventing adversaries from establishing links between addresses. We review mixing protocols and approaches adopted in privacy-native currencies.
@@
Different mixing protocols have been proposed, and we summarize them in Appendix A, Table LABEL:tab:sota-mixing. The naïve approach for mixing transactions consists in a centralized service that collects transactions, and publishes on the ledger a single large transaction with multiple inputs and outputs. This hides the information regarding which inputs flows towards which output, thus limiting (if not preventing) linkability between sender and receiver. The Mixcoin centralized protocol considers an accountable mixer, which offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Blindcoin (Valenta and Rowan, 2015) extends this protocol by considering blind signatures for the warrants; this hides the link between input and output addresses of a user even from the mixer, achieving full unlinkability; however, the mixer can still thief coins. Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. In BlindHub (Qin et al., 2023), mixing is performed within payment channel hubs, where the hub acts as tumbler and uses blind signatures to hide the customers’ addresses from itself. Obscuro (Tran et al., 2018) is a mixing protocol that utilizes Trusted Execution Environments (TEEs) and enables participants to verify the mixer’s integrity via remote attestation. It is specifically designed to defend against two key threats: the participation rejection attack, where malicious users try to manipulating mixing by withdrawing their deposits before the TEE reads them, and the blockchain forking attack, where adversaries present stale blocks to mislead the mixer. Numerous papers also investigate decentralized protocols. CoinJoin (Maxwell, 2013) is a technique that allows multiple users to combine their transactions into a single joint transaction; the author does not propose a fully fledged protocol, but only describes the mixing idea. CoinShuffle (Ruffing et al., 2014) allow to combine multiple users’ inputs and outputs in a single transaction but may compromise anonymity if transaction values differ or the anonymity set is small. Later, Ruffing et al. (Ruffing et al., 2016) proposed CoinShuffle++, which uses a more efficient mechanism for anonymity, i.e., Dining Cryptographers Networks (DC-nets) instead of mix-nets: while a mix-net requires sequential processing, hence a number of communication rounds linear in the number of participants, DC-nets enable to process mixing in parallel, hence it requires only a constant number of communication rounds. Xim (Bissias et al., 2014) is a multi-round protocol for anonymously finding mix partners based on advertisements placed on the blockchain, with the property that no outside party can identify participants that pair up. SecureCoin (Ibrahim, 2017) creates an aggregated temporary deposit address and requires the collaboration of involved peers to generate novel destination addresses and collectively redistribute coins. Leveraging Confidential Transactions (which nonetheless is available only in a soft fork of Bitcoin), ValueShuffle (Ruffing and Moreno-Sanchez, 2017) provides full privacy besides parties anonymity. Confidentiality enables mixing funds of different values without compromising unlinkability. CoinParty (Ziegeldorf et al., 2018) is a decentralized mixing protocol that distributes the mixing process across a group of semi-trusted parties. They use an MPC protocol to jointly perform sensitive operations (i.e., encryption, decryption, and address shuffling), ensuring that no single party can link input and output addresses. Threshold signatures is used to enable claiming funds upon agreement of a majority of mixing nodes. It also includes mechanisms to detect and exclude misbehaving participants, increasing the overall protocol reliability.
@@
Decentralized tumblers have also been designed using smart contracts on Ethereum. Möbius (Meiklejohn and Mercer, 2018) uses ring signatures and stealth addresses to obfuscate the senders and recipients associations (as Monero does (Noether et al., 2016)). First, senders furnish both their funds and stealth keys to the smart contract; then, each recipient creates a ring signature to withdraw their funds from the contract and transfer them to an ephemeral address. The size of the anonymity set is limited to the size of the ring, and the gas cost of the withdrawing transaction increases linearly with the size of the ring. MixEth (Seres et al., 2019) aims to minimize gas costs by utilizing Neff’s verifiable shuffle (Neff, 2001) and off-chain messages. Once the sender deposits coins, the MixEth contract enters rounds of shuffling and challenging, during which anyone can challenge the correctness of the preceding shuffle. When recipients believe that sufficient shuffling has been performed, they can withdraw their funds from the MixEth contract. We also mention AMR (Le and Gervais, 2021) which disrupts the traceability connection between coins deposited and withdrawn by the same user. Participants deposit a predetermined quantity of coins into a smart contract, which establishes a Merkle tree structure over the deposits. To withdraw coins, a zk-SNARK prove coins inclusion in the Merkle tree. AMR leverages lending platforms to provide users with the opportunity to accrue interest on their deposited assets.
@@
Although improving unlinkability, mixing protocols are still vulnerable to transaction graph analysis to some extent. For example, if the mixing process does not involve adequate randomization or if there are patterns in the amounts or timing of transactions, likely connections can be identified, even after several rounds of mixing. Also, if a participant does not properly anonymize the change, it can be traced back to the sender. To maximize privacy, participants need to combine mixing protocols with other privacy-preserving techniques and follow best practices (e.g., distributing coins across multiple addresses) to further obscure their transaction history and reduce the risk of being traced through transaction graph analysis. Further details on mixing solutions for Bitcoin and Ethereum, e.g., in (Arbabi et al., 2023) and on transaction graph analysis, e.g., in (Ron and Shamir, 2013; Fleder et al., 2015; Chan and Olmsted, 2017).
@@
Privacy-preserving cryptocurrencies leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not address the issue of transaction amounts, which are still visible on the ledger. Moreover, existing protocols leveraging updatable public keys require the receiver to scan the ledger for incoming payments, which adds complexity for the user. Ring signatures (e.g., (Van Saberhagen, 2013; Noether et al., 2016)) build upon the idea of mixing, which redistributes funds in a manner that obscures the origins (and destinations) of the coins. With ring signatures, a sender creates a ring with multiple decoy inputs, making it impossible to determine which input belongs to the spender. However, the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set, thus compromising privacy. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computational demanding (see Section 4.5).

### v2

@@ 4.3.3. Transaction Value Confidentiality
Two main approaches conceal transaction details: commitments (e.g., (Ben Sasson et al., 2014; Androulaki and Karame, 2014; Noether et al., 2016; Poelstra, 2016; Garman et al., 2017; Fauzi et al., 2019; Wüst et al., 2019; Gross et al., 2021; Guan et al., 2022; Wüst et al., 2022; Beer et al., 2024; Elfadul et al., 2024; Tomescu et al., 2022; Androulaki et al., 2020; Narula et al., 2018)) and public key encryption (e.g., (Kiayias et al., 2022; Sarencheh et al., 2025; Bünz et al., 2020; Ma et al., 2021)).
@@
Camenisch et al. (Camenisch et al., 2006) use commitments with verifiable encryption to hide the commitment’s opening. In 2013, Maxwell proposed confidential transactions141414https://elementsproject.org/features/confidential-transactions/investigation, which use Pedersen commitments to hide transaction amounts while ensuring correctness (also using range proofs). Monero adopted and extended this technique by combining ring signatures for sender anonymity, obtaining RingCT (Noether et al., 2016). Zerocash (Ben Sasson et al., 2014) uses ZKPs to achieve full transaction privacy: a cryptographic commitment is made to a new coin, including its value, owner, and serial number; then, a ZKP ensures transaction validity without exposing sensitive data. Commitment-based schemes are the most widely adopted to obtain confidential transaction value, especially because their homomorphic versions allow to combine commitments without leaking sensitive data. To prevent users from creating negative or infinite amounts, range proofs prove that a committed value is greater than zero and within a valid range.
@@
In PEReDi (Kiayias et al., 2022), senders and receivers update their accounts by submitting encrypted transactions to several maintainers who operate a distributed ledger. Threshold ElGamal encryption (Elgamal, 1985) requires a quorum of maintainer to cooperate to decrypt transaction details. Authorities can trace both transactions and identities. PARScoin (Sarencheh et al., 2025) extends this approach to introduce an offline payment protocol that does not require interaction between sender and receiver. In Solidus (Cecchetti et al., 2017), Publicly Verifiable Oblivious RAMs (PVORMs) enable banks to update the accounts of their clients without revealing exactly which accounts are being updated. PVORM provides a ZKP demonstrating that the updates are correct with respect to the transaction triggering them.
A few solutions implement privacy-preserving digital payments using smart contracts on public blockchains. To ensure confidentiality, transactions are encrypted with the sender’s public key and supplemented with a NIZK proof for public verifiability (e.g., (Ma et al., 2021; Bünz et al., 2020)). For example, Zether (Bünz et al., 2020) uses smart contracts to define a privacy-preserving tokens that can be exchanged between public keys associated with a confidential balance. They use ElGamal encryption, which has homomorphic properties, for hiding each account’s balance. Zether introduces the -Bullets (an enhancement of the Bulletproofs range ZKP) to efficiently prove statements over the encrypted transfer value and the new sender balance. However, Zether currently allows only one anonymous transfer per epoch (to prevent double spending), and execution costs (namely, gas costs on Ethereum) are very high (Almashaqbeh and Solomon, 2022). Similarly, DSC (Ma et al., 2021) optimizes privacy within Ethereum-style smart contract environments, but does not support anonymity: it uses ZKPs and homomorphic encryption to hide balances and transferred amounts. User balances are stored on the ledger in encrypted form.
@@ 4.3.4. Network Anonymity
Network anonymity concerns the protection of sender and receiver IP addresses, which could otherwise reveal real-world identities. Most of the existing approaches do not investigate (or even mention) the network anonymity feature. Few works mention it but consider this issue as out of scope of their investigation (e.g., (Cecchetti et al., 2017; Baudet et al., 2023; Tomescu et al., 2022; Wüst et al., 2022)). Monero (Noether et al., 2016) and Traceable Monero (Li et al., 2021) explicitly support network anonymity by using TOR (The Onion Router151515https://www.torproject.org/) and I2P (Invisible Internet Project161616https://geti2p.net/en/) to hide the real IP address (Shi et al., 2024). While TOR routes traffic through multiple relays to obscure the user’s IP address, I2P is a network optimized for anonymous peer-to-peer communication, using fully decentralized routing and encryption to enable different anonymous services. Moreover, Monero uses Dandelion++ (Fanti et al., 2018), an improvement of Dandelion (Bojja Venkatakrishnan et al., 2017), which prevents malicious attackers from linking transactions with their source IPs. Dandelion++ uses randomization mechanisms for message forwarding, which happens in two phases: stem (anonymization) and fluff (broadcast). In the stem phase, a transaction is first relayed privately across a small, randomly chosen subset of nodes; each node probabilistically decides whether to continue forwarding the transaction along the stem or to switch to the second phase. In the fluff phase, the node broadcasts to the entire network using traditional flood propagation protocol.
@@ 4.3.5. Linkability and Unlinkability
Unlinkability ensures that transactions cannot be correlated or traced back to the same user (other definitions in (Chaum, 1983; Amarasinghe et al., 2019; Zhang, 2023)). Unlinkability can refer to the sender, receiver, or both. A few works break this property into two facets, internal and external. For instance, CoinJoin (Maxwell, 2013) does not preserve internal unlinkability as each participant in the group that mixes transactions can link entities within the group.
Obtaining unlinkability in public ledgers is challenging as linking between addresses in crypto-asset systems stems from different patterns (e.g., (Ron and Shamir, 2013; Chan and Olmsted, 2017; Béres et al., 2021; Victor, 2020; Zhong and Mueen, 2024; Lu et al., 2022; Ermilov et al., 2017; Kalodner et al., 2020)). A non exhaustive list of patterns and heuristics to link addresses is the following.
@@
Address reuse: Using the same address for multiple transactions make it easy to link payments to the same user171717This works because address are pseudonymous but not private.;
@@
Dust attack analysis: Small amounts of crypto-assets are sent to many addresses to track their movements (similar to a phishing attack); if dust is later spent with other transactions, it links addresses together.
Existing countermeasures primarily aim to obscure input-output relations within transactions by using mixing protocols and preventing address reuse. Mixing protocols disrupt common-ownership heuristics and help mitigate change-address attacks. Address-changing mechanisms, such as generating a fresh sender or receiver address for each transaction, further reduce linkability. In addition, using network-level anonymity tools helps conceal user identities. Finally, the simplest defense against dust attacks is to ignore dust transactions, thereby preventing adversaries from linking addresses. A review of techniques for improving unlinkability follows.
@@
Different mixing protocols have been proposed (see Appendix A, Table LABEL:tab:sota-mixing). The naïve approach for mixing transactions consists in a centralized service that collects transactions, and publishes on the ledger a single large transaction with multiple inputs and outputs. This hides the information regarding which inputs flows towards which output, thus limiting (if not preventing) linkability between sender and receiver. The Mixcoin centralized protocol considers an accountable mixer, which offers a signed warranty that will enable his customers to unambiguously prove if the mix has misbehaved (Bonneau et al., 2014). Blindcoin (Valenta and Rowan, 2015) extends this protocol by considering blind signatures for the warrants; this hides the link between input and output addresses of a user even from the mixer, achieving full unlinkability. Tumblebit (Heilman et al., 2016) prevents coin thief using an on-chain escrow and ZKPs. In BlindHub (Qin et al., 2023), mixing is performed within payment channel hubs, where the hub acts as tumbler and uses blind signatures to hide the customers’ addresses from itself. Obscuro (Tran et al., 2018) is a mixing protocol that utilizes Trusted Execution Environments (TEEs) and enables participants to verify the mixer’s integrity via remote attestation.181818Obscuso is specifically designed to defend against two threats: the participation rejection attack, where malicious users try to manipulating mixing by withdrawing their deposits before the TEE reads them, and the blockchain forking attack, where adversaries present stale blocks to mislead the mixer. Numerous papers also investigate decentralized protocols. CoinJoin (Maxwell, 2013) is a technique that allows multiple users to combine their transactions into a single joint transaction; however, the author only describes the mixing idea. CoinShuffle (Ruffing et al., 2014) allow to combine multiple users’ inputs and outputs in a single transaction but may compromise anonymity if transaction values differ or the anonymity set is small. Later, Ruffing et al. (Ruffing et al., 2016) proposed CoinShuffle++, which uses a more efficient mechanism for anonymity, i.e., Dining Cryptographers Networks (DC-nets) instead of mix-nets: while a mix-net requires sequential processing, hence a number of communication rounds linear in the number of participants, DC-nets enable to process mixing in parallel, hence it requires only a constant number of communication rounds. Xim (Bissias et al., 2014) is a multi-round protocol for anonymously finding mix partners based on advertisements placed on the blockchain, with the property that no outside party can identify participants that pair up. SecureCoin (Ibrahim, 2017) creates an aggregated temporary deposit address and requires the collaboration of involved peers to generate novel destination addresses and collectively redistribute coins. Leveraging Confidential Transactions (which nonetheless is available only in a soft fork of Bitcoin), ValueShuffle (Ruffing and Moreno-Sanchez, 2017) provides full privacy besides parties anonymity. Confidentiality enables mixing funds of different values without compromising unlinkability. CoinParty (Ziegeldorf et al., 2018) is a decentralized mixing protocol that distributes the mixing process across a group of semi-trusted parties. They use an MPC protocol to jointly perform sensitive operations (i.e., encryption, decryption, and address shuffling), ensuring that no single party can link input and output addresses. Threshold signatures is used to enable claiming funds upon agreement of a majority of mixing nodes. It also includes mechanisms to detect and exclude misbehaving participants, increasing the overall protocol reliability.
@@
Decentralized tumblers have also been designed using smart contracts on Ethereum. Möbius (Meiklejohn and Mercer, 2018) uses ring signatures and stealth addresses to obfuscate the senders and recipients associations. First, senders furnish both their funds and stealth keys to the smart contract; then, each recipient creates a ring signature to withdraw their funds from the contract and transfer them to an ephemeral address. The size of the anonymity set is limited to the size of the ring, and the gas cost of the withdrawing transaction increases linearly with the size of the ring. MixEth (Seres et al., 2019) aims to minimize gas costs by utilizing Neff’s verifiable shuffle (Neff, 2001) and off-chain messages. Once the sender deposits coins, the MixEth contract enters rounds of shuffling and challenging, during which anyone can challenge the correctness of the preceding shuffle. When recipients believe that sufficient shuffling has been performed, they can withdraw their funds from the MixEth contract.
@@
Mixing protocols are still vulnerable to transaction graph analysis to some extent. For example, if the mixing process does not involve adequate randomization or if there are patterns in the amounts or timing of transactions, likely connections can be identified, even after several rounds of mixing. To maximize privacy, participants need to combine mixing protocols with other privacy-preserving techniques and follow best practices (e.g., distributing coins across multiple addresses) to reduce the risk of being traced through graph analysis. Further details on mixing solutions for Bitcoin and Ethereum, e.g., in (Arbabi et al., 2023), and on transaction graph analysis, e.g., in (Ron and Shamir, 2013; Fleder et al., 2015; Chan and Olmsted, 2017).
@@
Privacy-native solutions with public ledgers leverage a variety of strategies to improve unlinkability, focusing primarily on address-changing, transaction mixing, and use of ZKPs. Each of these techniques adds a layer of privacy, with trade-offs in terms of anonymity, computational complexity, and vulnerability to analysis. One-time destination addresses, stealth addresses (e.g., (Van Saberhagen, 2013; Noether et al., 2016)), or updatable public keys (e.g., (Papadoulis et al., 2024; Fauzi et al., 2019)) help obtain recipient unlinkability. While this ensures the recipient’s privacy, it does not hide transaction amounts. Protocols leveraging updatable public keys require the receiver to periodically scan the ledger, adding complexity. With ring signatures (e.g., (Noether et al., 2016; Van Saberhagen, 2013)), a sender creates a ring with multiple decoy inputs, making it impossible to determine which input belongs to the spender. However, the privacy protection is limited by the number of decoys in the ring, and sophisticated heuristics can sometimes reduce the anonymity set. For the most robust privacy, ZKPs offer a powerful solution (e.g., (Ben Sasson et al., 2014; Poelstra, 2016)). They allow transactions to prove their validity without revealing any details about the sender, receiver, or transaction amount. With this approach, transactions are recorded as cryptographic proofs (instead of sender-receiver addresses), ensuring that outputs cannot be linked to inputs. Among ZK-based approaches, we mention the commitments and nullifiers strategy (e.g., (Ben Sasson et al., 2014; Tomescu et al., 2022)): a coin contains a commitment to its serial number, and spending the coin involves publishing a nullifier, i.e., a pseudo-random value derived from the coin serial number. This ensures unlinkability to the original coin, while marking it as spent. However, designing a system with commitments and nullifiers is more complex than ring signatures or stealth addresses, and is also more computational demanding (see Section 4.5).